OpenStudy (anonymous):

f(x)=1/2x-6 what is an equation for f^1(x)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[f(x)=\frac{1}{2}x-6\] or \[f(x)=\frac}1}{2x-6}\]?

OpenStudy (anonymous):

<3 @misty1212

OpenStudy (misty1212):

lol i meant \[f(x)=\frac{1}{2x-6}\] i hope it is the first one !

OpenStudy (anonymous):

its the first one :)

OpenStudy (misty1212):

@DaWizjr \[\color\magenta\heartsuit\]

OpenStudy (misty1212):

\[f(x)=\frac{1}{2}x-6\] means a) divide by 2 b) subtract 6 inverse will be a) add 6 b) multiply by 2

OpenStudy (anonymous):

1/2( one half)x -6 , not the 1 over everything

OpenStudy (misty1212):

write it as \[f^{-1}(x)=2(x+6)\]

OpenStudy (misty1212):

or if you prefer \[f^{-1}(x)=2x+12\]

OpenStudy (misty1212):

like how you would solve \[\frac{1}{2}x-6=10\] for \(x\) add \(6\) multiply by \(2\)

OpenStudy (anonymous):

shoot I'm sorry